# On measures derived from orbital integrals

### v1

@@ Abstract.
Keywords. orbital integral, orbital measures, piecewise polynomial measure, apolar inner product, Fisher-Bombieri inner product, reproducing kernel.
@@ 1. Introduction
In this article we develop a novel approach to the derivation of some piecewise polynomial measures or Duistermat-Heckman type measures from Harish-Chandras (HC) orbital integral formula in the context of compact semi-simple Lie groups. These measures are derived from orbital measures, that is, invariant probability measures on adjoint orbits of compact semi-simple Lie groups. Specifically, we derive the pushforward of an orbital measure by the projection to the Cartan algebra, and we derive the radial part of the convolution of two orbital measures. These problems were previously addressed in [Z16, O13, F15] and in [F19, CMZ19, Z18]. They are probabilistic versions of the Schur-Horn theorem [H54] and Horn’s problem [H62]. Given hermitian matrices $X$ and $Y$ with fixed eigenvalues the Schur-Horn theorem characterizes the possible diagonal entries of $X$ as the convex closure of the permutation of eigenvalues, while Horn’s problem asks for a characterization of the possible eigenvalues of the sum $X+Y$. The possible eigenvalues are given by a set of linear inequalities and the problem was solved by Klyatchko [K98].
The approach in this article applies directly to compact semi-simple Lie groups and is based on operations which can be done in finite dimensional spaces of polynomials endowed with the apolar inner product, which is the Segal-Bargmann-Fock space inner product restricted to polynomials. This operation if performed to both sides of an equation where the Fourier-Laplace transform of the unknown measure is written as an exponential polynomial involving the discriminant. The framework in the article was motivated in part by [M25] where the Harish-Chandra-Itzykson-Zuber (HCIZ) integral formula was put in the context of Segal-Bargmann spaces, and also [PS09] where divided difference operators and their adjoints were used to solve algebraic problems. In this article we took finite degree parts of power and performed operations on these polynomials. Proofs in which finite degree terms are considered separately and an exponential function is constructed at the end are not uncommon in related literature, see for example the character expansion in the HCIZ integral in [MS17] and the proof of the Duistermaat-Heckman localization formula [MS17].
@@ 2. Preliminaries
We review the results which are necessary to prove the main results of the article, such as the Harish-Chandra integral formula and operators between spaces of polynomials endowed with the apolar inner product. Most results are known, to the best of our knowledge only Propositions 2.5, 2.2 and 2.9 are new.
@@ 2.1. Orbital measures and the Harish-Chandra integral formula
Let $G$ be a compact, semisimple and connected Lie group with Lie algebra $\operatorname{\mathfrak{g}}$. The Killing form is the $\operatorname{Ad}$-invariant inner product in $\operatorname{\mathfrak{g}}$ and is denoted with $\langle\cdot,\cdot\rangle$, $\operatorname{\mathcal{W}}$ is the Weyl group, $\operatorname{\epsilon}(w)$ is the sign of $w\in\operatorname{\mathcal{W}}$, that is $\operatorname{\epsilon}(w)=(-1)^{|w|}$ where $|w|$ is the number of reflections necessary to generate $w$. A compact Lie group admits a unique invariant probability measure called the Haar measure and the adjoint action induces a unique invariant probability measure $\nu_{a}$ on the orbits $\operatorname{\mathcal{O}}_{a}=\operatorname{Ad}_{G}(a)$ for $a\in G$. In [HC57] Harish-Chandra proved a formula for orbital integrals in Lie algebras in the case of compact, connected and semi-simple Lie groups. The orbital integral admits an expression as an exponential polynomial. It is usually written
$\Delta(x)\Delta(y)\int_{G}e^{\langle\operatorname{Ad}_{g}x,y\rangle}dg=\left[\Delta,\Delta\right]\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(x),y\rangle},$
@@
is the product of positive roots $\Phi^{+}$ considered as linear functionals. The term $\left[\Delta,\Delta\right]$ is computed with the apolar inner product which we a are going to introduce in the next section. We usually consider the element $a\in\operatorname{\mathfrak{t}}$ of the orbit $\operatorname{\mathcal{O}}_{a}$ as fixed and we write when $\Delta(a)\neq 0$
@@ 2.2. Spaces of polynomials and operators between them
for $f,g\in\operatorname{\mathcal{P}}^{k}(V)$. Here $f(\partial)$ is the polynomial $f(\frac{\partial}{\partial x_{1}},\dots,\frac{\partial}{\partial x_{n}})$. This inner product is sometimes called Fisher or Bombieri inner product. If $(x_{1},\dots,x_{n})$ are coordinates given by an orthonormal basis of $V$ then the monomials given by $x^{\beta}/\sqrt{\beta!}$ are orthonormal, where $\beta$ is multi-index and $\beta!:=\beta_{1}!\beta_{2}!\dots\beta_{n}!$. Note that
@@
$g^{*}(\partial)f|_{z=0}=\frac{1}{\pi^{n}}\int_{\operatorname{\mathbb{C}}^{n}}f(z)\overline{g(z)}e^{-|z|^{2}}dz$
@@
A multivariate version of Taylor’s theorem applied to polynomials asserts that for a polynomial $f$
$f(x+a)=e^{\langle a,\partial\rangle}f(x)=\sum_{\beta}a^{\beta}\partial^{\beta}f(x)$
If we denote by $T_{a}f(x)=f(x+a)$ the translation operator on polynomials then $T_{a}=e^{\langle a,\partial\rangle}$ as operators on polynomials. In the space $\operatorname{\mathcal{P}}^{k}(V)$ we define for $a\in V$ the functions
@@
which are truncations of exponentials $e^{\langle x,a\rangle}$.
@@
where we used the definition of the the apolar inner product and Taylor’s theorem for polynomials. It also follows from the reproducing property of Segal-Bargmann spaces restricted to polynomials, see Section $5$ in [Ha00]. ∎
Consider the case when the inner product space is endowed with a finite refection group $\operatorname{\mathcal{W}}$ which has an alternating character $\epsilon:\operatorname{\mathcal{W}}\to\{1,-1\}$. We can define the spaces
@@
Define the orthogonal projections $P_{\operatorname{alt}}$ and $P_{\operatorname{sym}}$ on $\operatorname{\mathcal{P}}^{k}(V)$ given by
$P_{\operatorname{alt}}f(x)=\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)f(w(x))\qquad\mbox{and}\qquad P_{\operatorname{sym}}f(x)=\sum_{w\in\operatorname{\mathcal{W}}}f(w(x)).$
@@
For $a\in V$ the function $r_{a}^{k}$ has the reproducing property $\left[r_{a}^{k},g\right]=g(a)$ in $\operatorname{\mathcal{P}}_{\operatorname{alt}}^{k}(V)$.
@@
where we used definition of $r_{a}^{k}$, $P_{\operatorname{alt}}^{*}=P_{\operatorname{alt}}$ and the reproducing property of $q_{a}^{k}$. ∎
@@
The creation and annihilation operator are fundamental in the theory of Segal-Bargman space. They consist of multiplication by a variable and differentiation by the same variable. In the present context we define for $a\in V$ the multiplication operator
@@
For a constant $c\neq 0$ and unit norm $\alpha\in V$ we have $D_{c\alpha}=\frac{1}{c}D_{\alpha}$. Also, by a linear change of variables in the definition of $I_{\alpha}$ we can assume that $\|\alpha\|=1$. Since an orthogonal change of coordinates leaves the equation unchanged we can prove the identity for the standard basis vector $\alpha=e_{1}$. We can also verify the identity for polynomials $x_{1}^{k}f(x_{2},\dots,x_{n})$ and $x_{1}^{l}g(x_{2},\dots,x_{n})$ since polynomials of this type span the spaces of polynomials. Since the operators $D_{\alpha}$ and $I_{\alpha}$ don’t on the factors $f$ and $g$ and since
@@
It is easy to very that $\left[x_{1}^{k-1},x_{1}^{l}\right]=\left[x_{1}^{k},\frac{1}{l+1}x_{1}^{l+1}\right]$. If $k\neq l+1$ then both side are zero and if $k=l+1$ then both sides are equal to $l!$.
Until the end of this section we consider vector spaces which are the Cartan algebras $\operatorname{\mathfrak{t}}$ of Lie algebras of compact semi-simple Lie groups. One of the essential properties of the discriminant $\Delta:\operatorname{\mathfrak{t}}\to\operatorname{\mathbb{R}}$ is that it skew with respect to the action of $\operatorname{\mathcal{W}}$, $\Delta(w(x))=\operatorname{\epsilon}(w)\Delta(x)$. This follows from the fact that if $\alpha$ is a simple root the reflection through the plane $\langle\alpha,x\rangle=0$ sends $\alpha\mapsto-\alpha$ and permutes the other positive roots. The next proposition is well known.
@@
For each reflection $w_{\alpha}\in\operatorname{\mathcal{W}}$ we have $f(w_{\alpha}(x))=-f(x)$ so $f$ vanishes on the hyperplane $\langle\alpha,x\rangle=0$ and the polynomial $\langle\alpha,x\rangle$ divides $f$. Since the Lie algebra is semi-simple its root system is reduced, that is, the only scalar multiple of a root $\alpha$ is $\alpha$. Therefore, the positive roots are relatively prime, since they only divide each other if they are scalar multiples. Since for each positive root $\langle\alpha,x\rangle=0$ divides $f$ the discriminant $\Delta$ divides the polynomial $f$
@@
$\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})=\Delta\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}}).$
@@
We can define $D_{\Delta}=D_{\alpha_{l}}\circ D_{\alpha_{l-1}}\circ\dots\circ D_{\alpha_{2}}\circ D_{\alpha_{1}}$ for an ordering $\alpha_{1},\alpha_{2},\dots,\alpha_{l}$ of the $l=|\Phi|$ positive roots. The order of composition
@@
is defined by $I_{\Delta}=I_{\alpha_{1}}\circ I_{\alpha_{2}}\circ\dots\circ I_{\alpha_{l-1}}\circ I_{\alpha_{l}}$ for an ordering $\alpha_{1},\alpha_{2},\dots\alpha$ of the $l=|\Phi|$ positive roots $\Phi^{+}$.
@@
ince the order of the composition $D_{\Delta}=D_{\alpha_{l}}\circ D_{\alpha_{l-1}}\circ\dots\circ D_{\alpha_{2}}\circ D_{\alpha_{1}}$ is irrelevant and since $I_{\alpha}=D_{\alpha}^{*}$ for all $\alpha\operatorname{\mathfrak{t}}$ the conclusion of the statement of the theorem follows. ∎
@@
We have $D_{\Delta}(\operatorname{\mathbb{R}}\Delta)=\operatorname{\mathbb{R}}\mathbf{1}$ and $D_{\Delta}(\{\operatorname{\mathbb{R}}\Delta\}^{\perp})=\{\operatorname{\mathbb{R}}\mathbf{1}\}^{\perp}$. Therefore $I_{\Delta}\mathbf{1}=D^{*}_{\Delta}\mathbf{1}=d\Delta$ for $d\in\operatorname{\mathbb{R}}$. We have
$1=\left[D_{\Delta}\Delta,\mathbf{1}\right]=\left[\Delta,I_{\Delta}\mathbf{1}\right]=\left[\Delta,d\Delta\right]=d\left[\Delta,\Delta\right].$

### v2

@@ Abstract.
Keywords. orbital integral, orbital measures, piecewise polynomial measure, apolar inner product, Fisher Bombieri inner product, reproducing kernel.
@@ 1. Introduction
In this article we develop a novel approach to the derivation of some piecewise polynomial measures or Duistermat-Heckman type measures from Harish-Chandras (HC) orbital integral formula in the context of compact semi-simple Lie groups. These measures are derived from orbital measures, that is, invariant probability measures on adjoint orbits of compact semi-simple Lie groups. Specifically, we derive the pushforward of an orbital measure by the projection to the Cartan algebra, and we derive the radial part of the convolution of two orbital measures. These problems were previously addressed in [Z16, O13, F15] and in [F19, CMZ19, Z18]. They are probabilistic versions of the Schur-Horn theorem [H54] and Horn’s problem [H62]. Given hermitian matrices $X$ and $Y$ with fixed eigenvalues the Schur-Horn theorem characterizes the possible diagonal entries of $X$ as the convex closure of the permutation of eigenvalues, while Horn’s problem asks for a characterization of the possible eigenvalues of the sum $X+Y$. The possible eigenvalues are given by a set of linear inequalities and the problem was solved by Klyatchko in [K98].
The approach applies directly to compact semi-simple Lie groups and is based on operations which can be done in finite dimensional spaces of polynomials endowed with the apolar inner product, which is the Segal-Bargmann-Fock space inner product restricted to polynomials. This operation if performed to both sides of an equation where the Fourier-Laplace transform of the unknown measure is written as an exponential polynomial involving the discriminant. The framework in the article was motivated in part by [M25] where the Harish-Chandra-Itzykson-Zuber (HCIZ) integral formula was put in the context of Segal-Bargmann spaces, and also [PS09] where divided difference operators and their adjoints were used to solve algebraic problems. In this article we took finite degree parts of power series and performed operations on these polynomials. Proofs in which finite degree terms are considered separately and an exponential function is constructed at the end are not uncommon in related literature, see for example the character expansion in the HCIZ integral in [MS17] and the proof of the Duistermaat-Heckman localization formula [MS17].
@@ 2. Preliminaries
We review the results which are necessary to prove the main results of the article, such as the Harish-Chandra integral formula and operators between spaces of polynomials endowed with the apolar inner product. Most results are known, to the best of our knowledge only Propositions 2.5, 2.7, 2.9 and 2.10 are new.
@@ 2.1. Orbital measures and the Harish-Chandra integral formula
Let $G$ be a compact, semisimple and connected Lie group with Lie algebra $\operatorname{\mathfrak{g}}$. The Killing form is the $\operatorname{Ad}$-invariant inner product in $\operatorname{\mathfrak{g}}$ and is denoted with $\langle\cdot,\cdot\rangle$, $\operatorname{\mathcal{W}}$ is the Weyl group, $\operatorname{\epsilon}(w)$ is the sign of $w\in\operatorname{\mathcal{W}}$, that is, $\operatorname{\epsilon}(w)=(-1)^{|w|}$ where $|w|$ is the number of reflections necessary to generate $w$. A compact Lie group admits a unique invariant probability measure called the Haar measure and the adjoint action induces a unique invariant probability measure $\nu_{a}$ on the orbits $\operatorname{\mathcal{O}}_{a}=\operatorname{Ad}_{G}(a)$ for $a\in\operatorname{\mathfrak{g}}$. In [HC57] Harish-Chandra proved a formula for orbital integrals in Lie algebras in the case of compact, connected and semi-simple Lie groups, see the expository article [M21] and the references therein. The orbital integral admits an expression as an exponential polynomial. It is usually written
$\Delta(x)\Delta(y)\int_{G}e^{\langle\operatorname{Ad}_{g}x,y\rangle}dg=\frac{\left[\Delta,\Delta\right]}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(x),y\rangle},$
@@
is the product of the positive roots $\Phi^{+}$ considered as linear functionals. The term $\left[\Delta,\Delta\right]$ is computed with the apolar inner product which we a are going to introduce in the next section. We usually consider the element $a\in\operatorname{\mathfrak{t}}$ of the orbit $\operatorname{\mathcal{O}}_{a}$ as fixed and we write when $\Delta(a)\neq 0$
@@ 2.2. Spaces of polynomials and operators between them
for $f,g\in\operatorname{\mathcal{P}}^{k}(V)$. Here $f(\partial)$ is the polynomial $f(\frac{\partial}{\partial x_{1}},\dots,\frac{\partial}{\partial x_{n}})$. This inner product is sometimes called Fisher or Bombieri inner product. If $(x_{1},\dots,x_{n})$ are coordinates given by an orthonormal basis of $V$ then the monomials given by $x^{\beta}/\sqrt{\beta!}$ are orthonormal, where $\beta\in\operatorname{\mathbb{N}}^{n}$ is a multi-index and $\beta!:=\beta_{1}!\beta_{2}!\dots\beta_{n}!$. Note that
@@
$g^{*}(\partial)f(z)|_{z=0}=\frac{1}{\pi^{n}}\int_{\operatorname{\mathbb{C}}^{n}}f(z)\overline{g(z)}e^{-|z|^{2}}dz$
@@
A multivariate version of Taylor’s theorem applied to polynomials asserts that for a polynomial $f$ and $x,a\in V$
$f(x+a)=e^{\langle a,\partial\rangle}f(x)=\sum_{\beta}a^{\beta}\partial^{\beta}f(x).$
If for $a\in V$ we denote by $T_{a}f(x)=f(x+a)$ the translation operator on polynomials then $T_{a}=e^{\langle a,\partial\rangle}$ as operators on polynomials. In the space $\operatorname{\mathcal{P}}^{k}(V)$ we define for $a\in V$ the functions
@@
which are truncations of exponentials $e^{\langle x,a\rangle}$. We denote with $\operatorname{ev}_{a}(f)=f(a)$ the evaluation functionals at $a\in V$.
@@
where we used the definition of the the apolar inner product and Taylor’s theorem for polynomials. This property also follows from the reproducing property of Segal-Bargmann spaces restricted to polynomials, see Section $5$ in [Ha00]. ∎
Consider the case when the inner product space is endowed with a finite refection group $\operatorname{\mathcal{W}}$ which has an alternating character $\epsilon:\operatorname{\mathcal{W}}\to\{1,-1\}$. We can define the space of alternating polynomials
@@
In $\operatorname{\mathcal{P}}^{k}(V)$ we define the orthogonal projection onto the alternating polynomials $P_{\operatorname{alt}}$ and the orthogonal projection onto the symmetric $P_{\operatorname{sym}}$, they are given by
$P_{\operatorname{alt}}f(x)=\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)f(w(x))\qquad\mbox{and}\qquad P_{\operatorname{sym}}f(x)=\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}f(w(x)).$
@@
For $a\in V$ the function $r_{a}^{k}$ has the reproducing property
$\left[r_{a}^{k},g\right]=g(a)$
in $\operatorname{\mathcal{P}}_{\operatorname{alt}}^{k}(V)$.
@@
where we used the definition of $r_{a}^{k}$, $P_{\operatorname{alt}}^{*}=P_{\operatorname{alt}}$ and the reproducing property of $q_{a}^{k}$ sated in Proposition 2.2. ∎
@@
The creation and annihilation operator are fundamental in the theory of Segal-Bargman space. They consist of multiplication by a variable and differentiation by the same variable. In the context of spaces of polynomials we define for $\alpha\in V$ the multiplication operator
@@
For a constant $c\neq 0$ and unit norm $\alpha\in V$ we have $D_{c\alpha}=\frac{1}{c}D_{\alpha}$. Also, by a linear change of variables in the definition of $I_{\alpha}$ we can assume that $\|\alpha\|=1$. Since an orthogonal change of coordinates leaves the equation unchanged we can prove the identity for the standard basis vector $\alpha=e_{1}$. We can also verify the identity for polynomials $x_{1}^{k}f(x_{2},\dots,x_{n})$ and $x_{1}^{l}g(x_{2},\dots,x_{n})$ since polynomials of this type span the spaces of polynomials. Since the operators $D_{e_{1}}$ and $I_{e_{1}}$ don’t act on the factors $f$ and $g$ and since
@@
It is easy to very that $\left[x_{1}^{k-1},x_{1}^{l}\right]=\left[x_{1}^{k},\frac{1}{l+1}x_{1}^{l+1}\right]$. If $k\neq l+1$ then both side are zero and if $k=l+1$ then both sides are equal to $l!$. ∎
Until the end of this section we consider vector spaces which are the Cartan algebras $\operatorname{\mathfrak{t}}$ of Lie algebras of compact semi-simple Lie groups. One of the essential properties of the discriminant $\Delta:\operatorname{\mathfrak{t}}\to\operatorname{\mathbb{R}}$ is that it skew with respect to the action of $\operatorname{\mathcal{W}}$, $\Delta(w(x))=\operatorname{\epsilon}(w)\Delta(x)$ for $w\in\operatorname{\mathcal{W}}$. This follows from the fact that if $\alpha$ is a simple root the reflection through the plane $\langle\alpha,x\rangle=0$ sends $\alpha\mapsto-\alpha$ and permutes the other positive roots. The next proposition is well known.
@@
For each reflection $w_{\alpha}\in\operatorname{\mathcal{W}}$ we have $f(w_{\alpha}(x))=-f(x)$ so $f$ vanishes on the hyperplane $\langle\alpha,x\rangle=0$ and the linear polynomial $\langle\alpha,x\rangle$ divides $f$. Since the Lie algebra is semi-simple its root system is reduced, that is, the only scalar multiple of a root $\alpha$ is $\alpha$. Therefore, the positive roots are relatively prime, since they only divide each other if they are scalar multiples. Since for each positive root the polynomial $\langle\alpha,x\rangle=0$ divides $f$ the discriminant $\Delta$ divides the polynomial $f$. ∎
@@
$\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})=\Delta.\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}}).$
@@
We can define $D_{\Delta}=D_{\alpha_{l}}\circ D_{\alpha_{l-1}}\circ\dots\circ D_{\alpha_{2}}\circ D_{\alpha_{1}}$ for an ordering $\alpha_{1},\alpha_{2},\dots,\alpha_{l}$ of the $l=|\Phi^{+}|$ positive roots. The order of the composition
@@
is defined by $I_{\Delta}=I_{\alpha_{1}}\circ I_{\alpha_{2}}\circ\dots\circ I_{\alpha_{l-1}}\circ I_{\alpha_{l}}$ for an ordering $\alpha_{1},\alpha_{2},\dots\alpha_{l}$ of the $l=|\Phi^{+}|$ positive roots $\Phi^{+}$.
@@
Since the order of the composition $D_{\Delta}=D_{\alpha_{l}}\circ D_{\alpha_{l-1}}\circ\dots\circ D_{\alpha_{2}}\circ D_{\alpha_{1}}$ is irrelevant and since $I_{\alpha}=D_{\alpha}^{*}$ for all $\alpha\in\operatorname{\mathfrak{t}}$ by Proposition 2.7 the conclusion of the statement of the theorem follows. ∎
@@
We have $D_{\Delta}(\operatorname{\mathbb{R}}\Delta)=\operatorname{\mathbb{R}}\mathbf{1}$ and $D_{\Delta}(\{\operatorname{\mathbb{R}}\Delta\}^{\perp})=\{\operatorname{\mathbb{R}}\mathbf{1}\}^{\perp}$. Therefore $I_{\Delta}\mathbf{1}=D^{*}_{\Delta}\mathbf{1}=d\Delta$ for a $d\in\operatorname{\mathbb{R}}$. We have
$1=\left[\mathbf{1},\mathbf{1}\right]=\left[D_{\Delta}\Delta,\mathbf{1}\right]=\left[\Delta,I_{\Delta}\mathbf{1}\right]=\left[\Delta,d\Delta\right]=d\left[\Delta,\Delta\right].$